OpenStudy (anonymous):

Need help! (Algebra 1)...

OpenStudy (lucaz):

?

OpenStudy (lucaz):

area=bh/2=47 b=x and h=x

OpenStudy (anonymous):

a=1/2 base*height 47=1/2*x*x 94=x^2 x=\[\sqrt{94}\] 9.7

OpenStudy (anonymous):

ok first The Area of a triangle is A= \[\frac{ 1 }{ 2 }*B*H\] Therefore \[47=\frac{ 1 }{ 2 }*X ^{2}\] Solve by Multiplying both sides by 2\[47*2=\frac{ 1 }{2 } X ^{2}*2\] \[94=X ^{2}\] Now solve for the square root \[\sqrt{94}=\sqrt{X ^{2}}\] \[X=9.7\]

OpenStudy (anonymous):

can you guys help me with this one?

OpenStudy (anonymous):

@kola908 thanks good explanation!

OpenStudy (anonymous):

you must consider the y intercept of the graph, the y intercept of the given graph is -1, a number which is lower than -1 is -3

OpenStudy (anonymous):

so im correct right?

OpenStudy (anonymous):

yes